OpenStudy (anonymous):

Medal for help :) x/4 = 12/6 A. x=6 B. x=12 C. x=8***My answer D. x=48

jimthompson5910 (jim_thompson5910):

Plug x = 8 into the original equation and show that please

OpenStudy (anonymous):

x/4 = 12/6 x = 12/6 (4) x = 12*4/6 x = 48/6 x = 8

OpenStudy (anonymous):

8 is the right answer

OpenStudy (jgirl128):

@JuniorEinstein1 great job, but next time you post a question, show your work so we can see you truly understand it :)

OpenStudy (jgirl128):

good job though!